OpenStudy (anonymous):

Find the maximum or minimum value of the function defined by f(x)=-x^2-6x-8

OpenStudy (anonymous):

the maximum or minimum value is the vertex of the function

myininaya (myininaya):

do -b/(2a) and then do f(-b/(2a))

OpenStudy (saifoo.khan):

the value of x^2 is in -ve so sad face.. Max point.

OpenStudy (saifoo.khan):

Of the value of x^2 is in +ve, then there will be a happy face.. Min Point,

OpenStudy (saifoo.khan):

Now judge yours!

OpenStudy (anonymous):

maximum of two

myininaya (myininaya):

f(-(-6)/2(-1))=f(6/(-2))=f(-3)=-(-3)^2-6(-3)-8

myininaya (myininaya):

i dont get 2

myininaya (myininaya):

i get 1 we can also do it another way but it is longer f(x)-(x^2+6x)-8=-(x^2+6x+9)-8+9 =-(x+3)^2+1 vertex is (-3,1) the max value is 1